OpenStudy (anonymous):

x+3=-(2x+2) step by step plz

OpenStudy (anonymous):

step 1: remove parentheses using distributive law \[x+3=-2x-2\]

OpenStudy (anonymous):

step 2: add 2x to both side \[x+2x+3=-2x+2x-2\] \[3x+3=-2\]

OpenStudy (anonymous):

step 3: subtract 3 from both sides \[3x+3-3=-2-3\] \[3x=-5\]

OpenStudy (anonymous):

step 3: divide by 3 \[\frac{3x}{3}=\frac{-5}{3}\] \[x=-\frac{5}{3}\]

OpenStudy (anonymous):

the answer is 1 3/2

OpenStudy (anonymous):

that is all the steps one by noe

OpenStudy (anonymous):

1 2/3

OpenStudy (anonymous):

not to the question solve for x \[x+3=-(2x+2)\] it isn't.

OpenStudy (anonymous):

if you want you can write \[-\frac{5}{3}=-1\tfrac{2}{3}\] but is it definitely negative